Given that sine theta sinθ= 13/85 and cosine theta <​0, determine the values of the sine and cosine functions for 2thetaθ.
1
Expert's answer
2019-04-05T09:57:47-0400

Find "cos\theta" first ("cos\theta < 0") :

"cos \theta = - \sqrt{1 - sin^2 \theta} = -\sqrt {1 - (13/85)^2} = -84/85"

Find "sin 2 \theta" :

"sin 2 \theta = 2 sin \theta cos \theta = - 2 \cdot \frac{13}{85} \cdot \frac {84}{85} = - \frac{2184}{7225}"

Find "cos 2 \theta" :

"cos 2 \theta = cos^2 \theta - sin^2 \theta = (-\frac{84}{85})^2 - (\frac{13}{85})^2 = \frac{6887}{7225}"